7. Find the slope of the line passing thru these points
a. ( -2, 4) and ( -4, 8)
b. (3 , 5) and ( -5, -5)

Answer by Alan3354(69443)

7. Find the slope of the line passing thru these points
Don't put spaces in the points.
a. (-2,4) and (-4,8)
Slope m = diffy/diffx
m = (8-4)/(-4 -(-2)) = 4/(-2)
m = -2
=============
b. (3 , 5) and ( -5, -5)
Do it like the 1st one.